OpenStudy (anonymous):

I need help figuring out a related rates problem. A ladder 13feet long is leaning against the side of a building. If the foot of the ladder is pulled away from the building at a constant rate of 2/3 feet per second, how fast is the area of the triangle formed by the ladder, the building and the ground changing (in feet squared per second) at the instant when the top of the ladder is 12 feet above the ground?

OpenStudy (anonymous):

got a formula for the area of the triangle?

OpenStudy (anonymous):

I have in mind first using Pythagorean formula to find the height of the triangle formed by the 13 feet long leaning against the side of the building.

OpenStudy (anonymous):

A = 1/2 b*h

OpenStudy (anonymous):

Since I don't have the height, I have to use pythagorean to find h, right?

OpenStudy (anonymous):

ok lets use \(x\) and \(y\) instead

OpenStudy (anonymous):

making \[A=\frac{1}{2}xy\] you are looking for \[A'\]

OpenStudy (anonymous):

you got a formula for \(A'\)?

OpenStudy (anonymous):

uhm not yet

OpenStudy (anonymous):

don't think too hard, just write it (product rule)

OpenStudy (anonymous):

I have to find the derivative of A=1/2xy or A=1/2x(sqrt144-x^2) ?

OpenStudy (anonymous):

oh no just \(A(t)=\frac{1}{2}x(t)y(t)\)

OpenStudy (anonymous):

or just \[A=\frac{1}{2}xy\] i guess you can do it the other way too, but it is going to be annoying i can show you why they are the same when we are done the easy way

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what is \(A'\) ?

OpenStudy (anonymous):

\[\frac{ dA }{ dt } =\frac{ 1 }{ 2 } y\frac{ dx }{ dt } + \frac{ 1 }{ 2 } x\frac{ dy }{ dt } \]

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

ups

OpenStudy (anonymous):

yes, i would have written \[A'=\frac{1}{2}(xy'+yx')\]same thing only quicker

OpenStudy (anonymous):

now you are almost done you know 3 out of those four numbers

OpenStudy (anonymous):

what is y though?

OpenStudy (anonymous):

well we didn't say you want to call y the base or the height?

OpenStudy (anonymous):

oh, well seeing it as b * h I suppose the height

OpenStudy (anonymous):

probably should have specified at the beginning but up until now all has been symmetric, so it didn't matter

OpenStudy (anonymous):

ok so we call \(y\) the height 

OpenStudy (anonymous):

then you are told the following the foot of the ladder is pulled away from the building at a constant rate of 2/3 feet per second makes \[x'=\frac{2}{3}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you are told \[y=12\]

OpenStudy (anonymous):

pythagoras tells you \(x=5\)

OpenStudy (anonymous):

oo I see

OpenStudy (anonymous):

the only thing missing is \(y'\)

OpenStudy (anonymous):

so so in the case of "and the ladder changing at the instant the bottom of the ladder is 12 feet from the wall?" I would use y= sqrt(144-x^2) right

OpenStudy (anonymous):

i wouldn't but you can

OpenStudy (anonymous):

then that 12 would just be x instead of y

OpenStudy (anonymous):

i would use \[x^2+y^2=13\\ 2xx'+2yy'=0\\ y'=-\frac{x}{y}x'\]

OpenStudy (anonymous):

no no it would be correct, what you wrote

OpenStudy (anonymous):

\[y= \sqrt{169-x^2}\] actually

OpenStudy (anonymous):

and therefore \[y'=-\frac{xx'}{\sqrt{169-x^2}}\]

OpenStudy (anonymous):

oh yes lol

OpenStudy (anonymous):

in any case that is exactly what i wrote

OpenStudy (anonymous):

\[y'=-\frac{x}{y}x'\]

OpenStudy (anonymous):

because the denominator is \(y\)

OpenStudy (anonymous):

now you are good to find \(y'\)

OpenStudy (anonymous):

What is y' here?

OpenStudy (anonymous):

I know x' is the constant rate

OpenStudy (anonymous):

and you have all the numbers you need for \[A'=\frac{1}{2}(xy'+yx')\]

OpenStudy (anonymous):

to find \(y'\) you can use this \[y'=-\frac{xx'}{\sqrt{169-x^2}}\] what you wrote with \(x=5\) and \(x'=\frac{2}{3}\) or you can use \[y'=-\frac{x}{y}x'\]with \(x=5,y=12,x'=\frac{2}{3}\)

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

Yes, I'm solving it now and see what I get as result.

OpenStudy (anonymous):

good luck it is pretty much arithmetic from here on in so be careful with it, that is where all the mistake are usually made at least on my end

OpenStudy (anonymous):

@satellite73 is the answer negative?

OpenStudy (anonymous):

@satellite73 also how would this be different if "the ladder changing at the instant the bottom of the ladder is 12 feet from the wall?" from "the instant when the top of the ladder is 12 feet above the ground?"

OpenStudy (anonymous):

nvm I just figured out the difference in between both.

OpenStudy (anonymous):

I still have the doubt though about the answer. If I solve it when the ladder at the bottom is 12 then I get a negative answer but when I solve it when the ladder at the top is 12 then I get a positive answer. Even though I am searching for the area of the triangle, why is it affected by both?

OpenStudy (anonymous):

I just figured out my answer is correct. Now I am just wondering why the area is negative when the bottom is 12 and positive when top is 12.

OpenStudy (anonymous):

or why it decreases when bottom is 12 and increases when top is 12.

OpenStudy (anonymous):